OpenStudy (anonymous):

Part A: Use the properties of exponents to explain why 16 raised to the power of 1 over 4 is called the fourth root of 16. (5 points) Part B: The length of a rectangle is 5 units and its width is square root of 5 unit. Is the area of the rectangle rational or irrational? Justify your answer. (5 points)

OpenStudy (anonymous):

I would really appreciate some help with this. I have trouble with radicals and if someone could explain this question to me and help me answer it I'd be really grateful!

OpenStudy (jdoe0001):

\(\bf a^{\frac{n}{m}} = \sqrt[m]{a^n}\qquad thus\\\quad \\ 16^{\frac{1}{4}}\implies \sqrt[4]{16}\\ \quad \\ 16\implies 2\times 2\times 2\times 2\implies 2^4\qquad thus\\ \quad \\ \sqrt[4]{16}\implies \sqrt[4]{2^4}\implies 2\) thus "2" is a number that when multiplied FOUR TIMES by itself is 16, thus the FOURTH ROOT

OpenStudy (anonymous):

Thank you so much!! Could you help me with Part B?

OpenStudy (jdoe0001):

 is it an irrational number? well, check for yourself --> http://www.mathsisfun.com/irrational-numbers.html